# Practical hybrid decoding scheme for parity-encoded spin systems

### v2

@@ V.2 Spin-flipping mechanism and decoding cost
This paper does not discuss decoding cost in any further detail. Let us recall that the purpose of this paper is to propose the BF decoding and demonstrate its potential as a practical post-processing algorithm for a pre-processing stochastic algorithm, such as a QA. Since the development of QA devices is ongoing, it is not easy to demonstrate our BF decoding algorithm as a post-processing step for an actual QA device. Instead, the potential of the BF decoding was demonstrated using a classical MCMC sampler as a pre-processor in this study. We believe that the present result depends principally on the properties of the BF decoding algorithm, not on the MCMC sampler. Furthermore, our result is consistent with the general belief that two decoding algorithms based on different mechanisms can be combined to solve problems that neither can solve alone. We believe that our BF decoding algorithm is promising for correcting readout errors of the QA devices due to measurement errors as well as dynamic errors that may be encountered during QA.
@@ VI Conclusions
This paper proposes a practical decoding algorithm to correct readout errors for the SLHZ model. Given the close connection between the SLHZ model and the classical LDPC codes, classical decoding techniques for LDPC codes are expected to help exploit the potential inherent in the SLHZ model. We proposed a simple BF algorithm based on iterative majority voting, a decoding algorithm for LDPC codes, which is much simpler than the standard BP algorithm. We demonstrated that the BF algorithm provides strong protection against i.i.d. noise for spin readouts of the SLHZ system and is as efficient as the BP algorithm. To study the tolerance to errors caused by a broader range of noise, we conducted a classical simulation of spin readouts on the SLHZ system using MCMC sampler. We found that BF decoding can correct leakage errors caused by dynamical and thermal excitations during MCMC sampling. Efficient decoding is possible if we reduce energy penalties, allowing thermal excitations to populate correctable leakage states more frequently. This is quite reasonable, given that stochastic sampling followed by classical error correction can be regarded as a two-stage hybrid decoding algorithm. Our observation suggests that the SLHZ model exhibits error correction against a broader range of noise models than the i.i.d. noise model when combined with post-readout BF decoding. Note that the BF algorithm does not suffer from the energy-barrier issue inherent to stochastic sampling. Thus, the two-stage hybrid decoding algorithm can be regarded as a practical implementation of the soft annealing proposed by Sourlas Sourlas (2005). Controlling the annealing parameters in the first-stage sampler is crucial for efficiently obtaining the correct state. Special attention should be paid to the fact that the optimal annealing parameters depend heavily on whether decoding is performed on the sampled readout.
@@
Since our demonstration used readouts sampled stochastically by a classical method, specifically MCMC sampling, we are not necessarily convinced that the BF decoding is valid for readouts obtained through an SLHZ-based QA device. Further research is needed to assess how effectively the QA device can collaborate with BF decoding under realistic conditions. Nevertheless, we believe that most of our insights stem from the intrinsic nature of BF decoding and are applicable regardless of the stochastic sampling mechanism employed in the first stage. For example, if measurement error is the primary source of readout error of the SLHZ system, our BF decoding algorithm offers a straightforward solution to mitigate it. In addition, it would be reasonable to expect that a two-stage hybrid computation combining QA and post-readout BF decoding may address issues neither method can solve independently. Our research also emphasizes the importance of selecting appropriate decoding algorithms to exploit the potential of SLHZ systems.
@@
In this study, we tested only a small number of instances $K_{14}$ to investigate the performance and characteristics of the BF decoding; however, this is insufficient to draw a general conclusion. Of particular importance for future work is investigating the size dependence of the SLHZ model on the performance. In our discussion, we suggested that post-readout BF decoding may partially compensate for the poor spin-update properties of the SLHZ system. In particular, it will be interesting to see how this compensation works well for a problem of arbitrary size. Albash et al. also pointed out another significant limitation of the SLHZ system: the strength of the energy penalty must grow with the problem size Albash et al. (2016). Introducing post-readout BF decoding may alleviate this limitation since it reduces the penalty strength. Both these issues are important for achieving scalability. More research is needed to determine the efficacy of post-readout decoding. On the other hand, even if the performance evaluation using states generated by PT and SQA were valid, questions remain regarding the performance evaluation based on the number of sweeps required to generate them. This is because, although MC simulation describes spin dynamics as a sequential update of spin states at each sweep — a serial, discrete-time evolution —, QA dynamics are quite different. Namely, the latter involves the parallel, continuous evolution of spins over time. Therefore, the number of MC sweeps required and QA performance are not necessarily correlated. It would be important to make a reasonable performance comparison when developing an actual QA device.
@@ Acknowledgements.
In Ref.Albash et al. (2016), the authors present minimum weight decoding (MWD), which aims to find an error pattern $\boldsymbol{e}$ with minimum Hamming distance from the hard-decided readouts $\boldsymbol{r}$ and with the same syndrome vector as $\boldsymbol{s}^{(4)}(\boldsymbol{r})$. They directly estimated the optimal error pattern $\boldsymbol{e}^{*}\in\{\pm 1\}^{N_{v}}$ from the global ground state of the following Hamiltonian:
@@
We recognize that the performance boost of the MCMC-BF hybrid decoding at a vanishing penalty strength provides evidence supporting the inherent error-correcting capability of the SLHZ model. Our study suggested that controlling the penalty strength is crucial to exploiting the potential of the SLHZ model. Nontheless, we do not claim that this performance boost is sufficient for the SLHZ model to outperform the ME model. We rather consider that it is at least essential for the SLHZ system to be competitive with the ME model. For supplementary information, we have confirmed in our preliminary experiments that ME model shows no performance boost after post-readout MVD.

### v3

@@ V.2 Spin-flipping mechanism and decoding cost
This paper does not discuss decoding cost in any further detail. Let us recall that the purpose of this paper is to propose the BF decoding and demonstrate its potential as a practical post-processing algorithm for a pre-processing stochastic algorithm, such as a QA. Since the development of QA devices is ongoing, it is not easy to demonstrate our BF decoding algorithm as a post-processing step for an actual QA device. Instead, the potential of the BF decoding was demonstrated using a classical MCMC sampler as a pre-processor in this study. We believe that the present result depends primarily on the high compatibility between the SLHZ model and the BF decoding algorithm, rather than on the MCMC sampler. Furthermore, our result is consistent with the general belief that two decoding algorithms based on different mechanisms can be combined to solve problems that neither can solve alone. We believe that our BF decoding algorithm is promising for correcting readout errors of the QA devices due to measurement errors as well as dynamic errors that may be encountered during QA.
@@ VI Conclusions
This paper proposed a practical decoding algorithm to correct readout errors tailored for the SLHZ model. Given the close connection between the SLHZ model and classical LDPC codes, classical decoding techniques for LDPC codes are expected to help exploit the latent potential of the SLHZ model. We proposed a simple BF algorithm based on iterative majority voting, a decoding algorithm for LDPC codes, which is much simpler than the standard BP algorithm. We demonstrated that the BF algorithm provides strong protection against i.i.d. noise for spin readouts of the SLHZ system and is as efficient as the BP algorithm. To study the tolerance to errors caused by a broader range of noise, we conducted a classical simulation of spin readouts on the SLHZ system using MCMC sampler. We found that BF decoding can correct leakage errors caused by dynamical and thermal excitations during MCMC sampling. Efficient decoding is possible if we reduce energy penalties, allowing thermal excitations to populate correctable leakage states more frequently. This is quite reasonable, given that stochastic sampling followed by classical error correction can be regarded as a two-stage hybrid decoding algorithm. Our observation suggests that the SLHZ model exhibits error correction against a broader range of noise models than the i.i.d. noise model when combined with post-readout BF decoding. Note that the BF algorithm does not suffer from the energy-barrier issue inherent to stochastic sampling. Thus, the two-stage hybrid decoding algorithm can be regarded as a practical implementation of the soft-annealing proposed by Sourlas [8]. Controlling the Lagrange weights in the first-stage sampler is crucial for efficiently obtaining the correct state. Special attention should be paid to the fact that the optimal Lagrange weights depend heavily on whether post-readout decoding is performed.
@@
Since our demonstration used readouts sampled stochastically by a classical method, specifically MCMC sampling, we are not necessarily convinced that the BF decoding is valid for readouts obtained through an SLHZ-based QA device. Further research is needed to assess how effectively the QA device can collaborate with BF decoding under realistic conditions. Nevertheless, we believe that most of our insights stem from the intrinsic nature of SLHZ model and BF decoding, and are applicable regardless of the stochastic sampling mechanism employed in the first stage. For example, if measurement error is the primary source of readout error of the SLHZ system, our BF decoding algorithm offers a straightforward solution to mitigate it. In addition, it would be reasonable to expect that a two-stage hybrid computation combining QA and post-readout BF decoding may address issues neither method can solve independently. Our research also emphasizes the importance of selecting appropriate decoding algorithms to exploit the potential of SLHZ systems.
@@
In this study, we tested only a small number of instances $K_{14}$ to investigate the performance and characteristics of the BF decoding; however, this is insufficient to draw a general conclusion. Of particular importance for future study is investigating the size dependence of the SLHZ model on the performance. In our discussion, we suggested that post-readout BF decoding may partially compensate for the poor spin-update properties of the SLHZ system. See the Appendix A in detail. In particular, it will be interesting to see how this compensation works well for a problem of arbitrary size. Albash et al. also pointed out another significant limitation of the SLHZ system: the strength of the energy penalty must grow with the problem size [24]. Introducing post-readout BF decoding may alleviate this limitation by reducing the penalty strength. Both these issues are important for achieving scalability. More research is needed to determine the efficacy of post-readout decoding.
@@ Acknowledgements.
In Ref.[24], the authors present minimum weight decoding (MWD), which aims to find an error pattern $\boldsymbol{e}$ with minimum Hamming distance from the hard-decided readouts $\boldsymbol{r}$ and with the same syndrome vector as $\boldsymbol{s}^{(4)}(\boldsymbol{r})$. They directly estimated the optimal error pattern $\boldsymbol{e}^{*}\in\{\pm 1\}^{N_{v}}$ from the global ground state of the following Hamiltonian:
@@
We recognize that the performance boost of the MCMC-BF hybrid decoding at a vanishing penalty strength provides evidence supporting the inherent error-correcting capability of the SLHZ model. Our study suggested that controlling the penalty strength is crucial to exploiting the potential of the SLHZ model. Nonetheless, we do not claim that this performance boost is sufficient for the SLHZ model to outperform the ME model. We rather consider it essential that the SLHZ system be competitive with the ME model. For supplementary information, we have confirmed in our preliminary experiments that ME model shows no performance boost after post-readout MVD.
Appendix B Supplementary result on experimental demonstration
In section IV, we demonstrated that the post-readout BF decoding algorithm can eliminate not only i.i.d. errors but also leakage errors due to dynamic and thermal excitations during the MCMC sampling process. However, when we imagine that the BF decoding is applied to the readout of actual QA devices rather than MCMC simulation, several types of errors may occur simultaneously. In this appendix, we assume i.i.d. errors in the readouts of MCMC sampler which might be a realistic model of readout errors for imperfect annealing devices, and investigate their impact on the performance of the MCMC-BF hybrid decoding. We present a result showing that the MCMC-BF decoder is robust to such readout errors to some extent.
Figure 12: Success probability for MCMC-BF decoder as a function of bit error rate $\varepsilon$ for the readouts of the MCMC sampler with four sample sizes. The sample size is expressed in a unit $N_{s}=\tbinom{K}{2}$, where $K=14$.
We simulate noisy readout $\hat{\boldsymbol{r}}^{\prime}$ by multiplying the error matrix $\hat{\boldsymbol{e}}$ generated by a pseudo-random number generator and the readout matrix $\hat{\boldsymbol{r}}$, i.e., $\hat{\boldsymbol{r}}^{\prime}=\hat{\boldsymbol{r}}\circ\hat{\boldsymbol{e}}$ , where $\hat{\boldsymbol{r}}$ was obtained from the same spin glass problem ($K_{14}$) as the main text. Subsequently, we applied the BF decoding to $\hat{\boldsymbol{r}}^{\prime}$, and investigated how the performance of the MCMC-BF hybrid decoder changes. To perform a statistical evaluation, we ran 1000 independent experiments. Figure12 shows the success probabilities for obtaining correct state when incorporating i.i.d. errors into the readouts of the MCMC sampler with four sample sizes. In this figure, we assumed that each spin flips with the bit error rate $\varepsilon$ $(0\leq\varepsilon\leq 0.5)$, and that the Lagrange weights are the values where the MCMC-BF hybrid decoder achieves maximum efficiency (corresponding to arrow B in Fig.9). As expected, the probabilities reduce as $\varepsilon$ increases and reach almost zero at $\varepsilon=0.5$. It was confirmed that decoding succeeds even when $\varepsilon\neq 0$, and, in particular, that the probability remains nearly unchanged in the region where $\varepsilon\leq 0.1$. This result suggests that the post-readout BF decoder can accommodate realistic noise in imperfect annealing devices.